OpenStudy (anonymous):

I did the work. I just need help putting it in standard form. Derive the equation of the parabola with a focus at (4, −7) and a directrix of y = −15. Put the equation in standard form. My answer in vertex form: 1/16(x-4)^2-7

OpenStudy (phi):

your equation has the vertex at 4, -7 But you want a parabola that has its focus at 4,-7 I think the vertex is half way between the focus and the directrix.

OpenStudy (anonymous):

My vertex was (4,-11).

OpenStudy (jdoe0001):

\(\bf \cfrac{1}{16}(x-\color{red}{4})^2\color{red}{-7}\) vertex

OpenStudy (anonymous):

Focus: (4,-7) Directrix: y= -15 Vertex: (4,-11) Distance: 4

OpenStudy (jdoe0001):

well, that's now what yours shows

OpenStudy (phi):

yes. so you want -11 in your equation not -7

OpenStudy (jdoe0001):

that's not rather

OpenStudy (anonymous):

Here are my choices: f(x) = x2 − 8x + 11 f(x) = x2 − 8x − 10 f(x) = x2 −1/2 x + 11 f(x) = x2 −1/2 x − 10 Right now, I think the answer is the first one.

OpenStudy (phi):

you should start with My answer in vertex form: 1/16(x-4)^2- 11

OpenStudy (anonymous):

Okay, how do I get from that to standard form?

OpenStudy (phi):

multiply out (x-4)(x-4)

OpenStudy (anonymous):

1/16[(x-4)(x-4)]-11

OpenStudy (phi):

multiply (x-4)(x-4) to get x^2 -4x -4x + 16 = x^2 -8x +16 so you have \[ \frac{ x^2}{16} - \frac{8}{16}x + \frac{16}{16} - 11 \] that simplifies to \[ \frac{ x^2}{16} - \frac{1}{2}x + 10\]

OpenStudy (phi):

*\[\frac{ x^2}{16} - \frac{1}{2}x - 10 \]

OpenStudy (phi):

all of your choices left out the divide by 16 on x^2

OpenStudy (anonymous):

Okay, Thanks so much!